OpenStudy (anonymous):

What is the simplified expression for the expression below? 7(x-4)-3(x+5)

OpenStudy (anonymous):

7x-28 - 3x -15 7x-3x -28-15 4x - 43

OpenStudy (anonymous):

What is the simplified expression for the expression below? 4(3x – 2) + 6x(2 – 1)

OpenStudy (anonymous):

12x - 8 + 12x -6 12x + 12x -8 -6 24x - 14

OpenStudy (anonymous):

Bernard solved the equation 5x+(-4)=6x+4 using algebra tiles. Which explains why Bernard added 5 negative x-tiles to both sides in the first step of the solution?

OpenStudy (anonymous):

hm.. well, you need to get the like numbers together. So 6x and 5x need to be on the same side. he subtracted it from both sides to cancel it out on one side and move it to the other.

OpenStudy (anonymous):

e wanted to create zero pairs on the left side of the equation to get a positive coefficient of x on the right side. He wanted to combine more negative tiles with the 4 negative tiles on the left side of the equation. He wanted to remove the larger coefficient on the left side of the equation. He wanted to have 10 x-tiles on the left side of the equation.

OpenStudy (anonymous):

a b c or d

OpenStudy (anonymous):

A because you can't have a negative x. Well, you can but that requires more work.

OpenStudy (anonymous):

In a basketball tournament, team A scored 6 more points than 3 times as many points as team B scored. Team C scored 45 more points than team B scored. The combined scores for all three teams was 476 points. Which statement about the number of points is true?

OpenStudy (anonymous):

Team A scored 85 points. Team A scored 255 points. Team C scored 261 points. Team C scored 130 point

OpenStudy (anonymous):

Well, B has the least amount of points. C has the second, and A has the most (I think). So A is out of the question as well as C. so it's B or D

OpenStudy (anonymous):

To solve the equation, Lorie applies the distributive property, combines like terms, then applies the addition and subtraction properties of equality to isolate the variable term on one side of the equation and the constant term on the other side. What are the possible coefficients of x after Lorie has completed these steps? 10(1/2x+2)-5=3(x-6)+1

OpenStudy (anonymous):

yeah, I'm not getting a solid solution. sorry.

OpenStudy (anonymous):

–32 and 2 –2 and 32 –2 and 2 –32 and 32 theres the option

OpenStudy (anonymous):

I think A. not positive though.